OpenStudy (anonymous):

(3g-4) + (4g+10) =90

OpenStudy (anonymous):

g=12.

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

3g - 4 + 4g + 10 = 90 Combine like terms 7g + 6 = 90 -6 -6 7g = 84 /7 /7 g = 12

OpenStudy (anonymous):

\[7g + 6 = 90\] \[7g = 84\] \[\text{Divide both sides by 7 to get}\] \[g = 12\]

OpenStudy (anonymous):

thank you!